# VQ-Style: Disentangling Style and Content in Motion with Residual Quantized Representations

### v1

@@ 4.7.2 Content Error
As [MSK22] train a locomotion controller rather than a style transfer model, their content deviation error cannot be measured. On the 100STYLE dataset, we find that much of the deviation arises from errors in predicted velocities during sharp turns in the content motion. Since the final motion is calculated by integrating the velocities over time, these velocity errors accumulate and lead to larger deviations toward the end of the sequence. Operating in a global reference space could potentially improve content preservation; however, we leave this investigation for future work.
@@ 4.8.2 Choice of swapping cut-off
Figure 13: Ablation on different cut-offs for style-content codebooks in style transfer application.
@@ 5 Conclusion and future work
Our method builds on the intuitive interpretation of content and style as coarse features versus finer details. However, formally defining and isolating these two components remains challenging. Even across existing public datasets, the criteria for what is considered style are inconsistent. For instance, kicking is annotated as content in the Aberman dataset, but is classified as a style in the 100STYLE dataset. Prior work has adopted various ways to define the style, e.g. via statistical features such as Gram matrices [HSK16], by shift and amplitude of instance normalization in latent space [AWL20], or by separating the non-temporal features through attention masks [TZCvdP22]. In our work, we view the content as coarse global motion structures, and style as the finer-grained local details. For a locomotion-focused dataset such as 100STYLES, global root trajectory and speed align more naturally with the content, and foot patterns, arm motions and other nuanced features align more closely with style. Similar to prior studies the boundary between content and style remains ambigious, and in practice the annotated style labels majorly influence the style-content distinction.
Furthermore, developing better quantitative metrics for style transfer remains an open task. For example, root trajectory error, a measure of content preservation, can be misleading. When applying a style like Drunk, the ideal output should intentionally alter the original motion’s root trajectory, making a low error score a poor indicator of a successful transfer. One can find similar counter examples for other metrics provided in the literature such as geodesic distance [GMZ24]. This is specifically more challenging for non-locomotion movements in which the expected result of style transfer might be ambiguous.
@@ Appendix A Background
VQ-VAEs [VDOV17] are powerful and compact representation models that have demonstrated success in encoding complex data, including human motion [GZWC22]. However, preserving high quality reconstruction when encoding a large and diverse dataset typically requires the codebook size to grow exponentially, posing challenges for scalability[S59, LKK22]. Residual VQ-VAE (RVQ-VAE) addresses this limitation by employing a hierarchical set of codebooks that encode residual errors at multiple levels[LKK22]. Rather than mapping the entire continuous latent vector to a single code drawn from a large codebook, the RVQ mechanism performs sequential refinement using $N$ codebooks $\mathcal{B}_{0},\mathcal{B}_{1},\cdots,\mathcal{B}_{N-1}]$. The first codebook quantizes the initial continuous latent vector, after which a residual is computed by subtracting the quantized output from the original latent. The second codebook then quantizes this residual, producing an additional refinement. This hierarchical residual quantization process continues through all $N$ codebooks, and the final latent representation is obtained by summing the quantized outputs from all stages.
@@
Rather than optimizing this objective directly, we employ an Exponential Moving Average (EMA) update which has been shown to improve the robustness and stability of codebook training [ŁCS20],
@@ Appendix B Training hyperparameters
The hyperparameters used for training our RVQ-VAE over different datasets are follows:
Table 4: Training hyperparameters values for reconstruction on 100Styles[MSK22] dataset. Values in parenthesis represent configuration used for training on Aberman[AWL20] dataset.
@@ Appendix C Metrics
We define style accuracy, $A_{S}$, as the accuracy of a style classifier on motion generated after applying style transfer. We train one classifier with identical architecture as [GMZ24]. This consists of four 1D convolutional layers followed by three deconvolutional layers and one linear layer. The classifiers are trained on all styles in their respective datasets and have test accuracies of 96.87% over the Aberman [AWL20] dataset and 98.57% over the 100STYLE dataset[MSK22].
@@ Appendix D Dataset details
We report $A_{S}$ & $D_{C}$ for train, test, and unseen styles subsets of the datasets. The train subset contains the same clips as the model was trained on. The test subset contains clips withheld during training but with the same styles seen during training.
For the 100STYLE dataset, the unseen dataset contains clips with styles that have never been seen by the model before. These styles are selected by choosing the last 10 styles in the alphabetical order. For comparison with LPN-Style [MSK22], we retrain our model on the same data split as theirs, where the unseen styles include HandsInPockets, Roadrunner, Skip, Star, WildArms and the model is trained on the remaining 95 styles.
For the Aberman dataset, we split the Aberman into train and test subsets similar to [GMZ24]. However, we withhold the complete Xia dataset as the unseen subset. Note that some of the styles in the Xia dataset are also present in the Aberman dataset. We choose to split this the datasets as such to for consistent comparison with [GMZ24]. However, it’s worth noting that the original GenMoStyle only uses Xia dataset as the content, never testing on unseen styles.
@@ Appendix E Per-style class content deviation
Figure 14: Plotting per class content deviation for train and test subsets of the 100Styles[MSK22] dataset.
Figure 14 plots the content deviation, $D_{C}$, for the train and test subsets of the 100STYLE[MSK22] dataset. We observe that the distribution of deviations is skewed, with most styles demonstrating a content deviation well below the mean deviation of their respective subsets. Specifically, of the 90 styles used for training, only 23 styles have an average deviation that exceeds the mean of the training subset. For the test subset, 25 styles have an average deviation greater than the test subset’s mean. In both training and testing, more energetic styles such as Rushed, Spin, Bounce, and Hop have worse content deviation.

### v2

@@ 4.7.2 Content Error
As [18] train a locomotion controller rather than a style transfer model, their content deviation error cannot be measured. On the 100STYLE dataset, we find that much of the deviation arises from errors in predicted velocities during sharp turns in the content motion. Since the final motion is calculated by integrating the velocities over time, these velocity errors accumulate and lead to larger deviations toward the end of the sequence. Operating in a global reference space could potentially improve content preservation; however, we leave this investigation for future work.
@@ 4.8.2 Choice of swapping cut-off
Figure 13: Ablation on different cut-offs for style-content codebooks in style transfer application. Left (pink): style clip, Middle: style transfer result, Right (blue): content clip.
@@ 5 Conclusion and future work
Our method builds on the intuitive interpretation of content and style as coarse features versus finer details. However, formally defining and isolating these two components remains challenging. Even across existing public datasets, the criteria for what is considered style are inconsistent. For instance, kicking is annotated as content in the Aberman dataset, but is classified as a style in the 100STYLE dataset. Prior work has adopted various ways to define the style, e.g. via statistical features such as Gram matrices [11], by shift and amplitude of instance normalization in latent space [1], or by separating the non-temporal features through attention masks [28]. In our work, we view the content as coarse global motion structures, and style as the finer-grained local details. For a locomotion-focused dataset such as 100STYLES, global root trajectory and speed align more naturally with the content, and foot patterns, arm motions and other nuanced features align more closely with style. Similar to prior studies the boundary between content and style remains ambigious, and in practice the annotated style labels majorly influence the style-content distinction.
Furthermore, developing better quantitative metrics for style transfer remains an open task. For example, root trajectory error, a measure of content preservation, can be misleading. When applying a style like Drunk, the ideal output should intentionally alter the original motion’s root trajectory, making a low error score a poor indicator of a successful transfer. One can find similar counter examples for other metrics provided in the literature such as geodesic distance [7]. This is specifically more challenging for non-locomotion movements in which the expected result of style transfer might be ambiguous.
@@
Acknowledgment
The authors would like to thank Dominik Borer for his technical support and Violaine Fayolle and Doriano van Essen for their artistic support during this project.
@@ Appendix A Background
VQ-VAEs [31] are powerful and compact representation models that have demonstrated success in encoding complex data, including human motion [8]. However, preserving high quality reconstruction when encoding a large and diverse dataset typically requires the codebook size to grow exponentially, posing challenges for scalability[23, 16]. Residual VQ-VAE (RVQ-VAE) addresses this limitation by employing a hierarchical set of codebooks that encode residual errors at multiple levels[16]. Rather than mapping the entire continuous latent vector to a single code drawn from a large codebook, the RVQ mechanism performs sequential refinement using $N$ codebooks $\mathcal{B}_{0},\mathcal{B}_{1},\cdots,\mathcal{B}_{N-1}]$. The first codebook quantizes the initial continuous latent vector, after which a residual is computed by subtracting the quantized output from the original latent. The second codebook then quantizes this residual, producing an additional refinement. This hierarchical residual quantization process continues through all $N$ codebooks, and the final latent representation is obtained by summing the quantized outputs from all stages.
@@
Rather than optimizing this objective directly, we employ an Exponential Moving Average (EMA) update which has been shown to improve the robustness and stability of codebook training [15],
@@ Appendix B Training hyperparameters
The hyperparameters used for training our RVQ-VAE over different datasets are detailed in Table 4.
Table 4: Training hyperparameters values for reconstruction on 100Styles[18] dataset. Values in parenthesis represent configuration used for training on Aberman[1] dataset.
@@ Appendix C Metrics
We define style accuracy, $A_{S}$, as the accuracy of a style classifier on motion generated after applying style transfer. We train one classifier with identical architecture as [7]. This consists of four 1D convolutional layers followed by three deconvolutional layers and one linear layer. The classifiers are trained on all styles in their respective datasets and have test accuracies of 96.87% over the Aberman [1] dataset and 98.57% over the 100STYLE dataset[18].
@@ Appendix D Dataset details
We report $A_{S}$ and $D_{C}$ for train, test, and unseen styles subsets of the datasets. The train subset contains the same clips as the model was trained on. The test subset contains clips withheld during training but with the same styles seen during training.
For the 100STYLE dataset, the unseen dataset contains clips with styles that have never been seen by the model before. These styles are selected by choosing the last 10 styles in the alphabetical order. For comparison with LPN-Style [18], we retrain our model on the same data split as theirs, where the unseen styles include HandsInPockets, Roadrunner, Skip, Star, WildArms and the model is trained on the remaining 95 styles.
For the Aberman dataset, we split the Aberman into train and test subsets similar to [7]. However, we withhold the complete Xia dataset as the unseen subset. Note that some of the styles in the Xia dataset are also present in the Aberman dataset. We choose to split this the datasets as such to for consistent comparison with [7]. However, it’s worth noting that the original GenMoStyle only uses Xia dataset as the content, never testing on unseen styles.
@@ Appendix E Per-style class content deviation
Figure 14: Plotting per class content deviation for train and test subsets of the 100Styles[18] dataset.
Figure 14 plots the content deviation, $D_{C}$, for the train and test subsets of the 100STYLE[18] dataset. We observe that the distribution of deviations is skewed, with most styles demonstrating a content deviation well below the mean deviation of their respective subsets. Specifically, of the 90 styles used for training, only 23 styles have an average deviation that exceeds the mean of the training subset. For the test subset, 25 styles have an average deviation greater than the test subset’s mean. In both training and testing, more energetic styles such as Rushed, Spin, Bounce, and Hop have worse content deviation.